Two charges lie on positive x-axis. Charge A (2.0 x 10-9 C) is 2.0 cm from the origin and Charge B is 4.0 cm from the origin. (-3.0 x 10-9 C). What is the total force exerted by these two charges on Charge C (5.0 x 10-9 C) located at the origin?     


1
Expert's answer
2022-02-09T13:42:30-0500

"F_1=\frac{kq_1q_2}{r^2}"


"F_1=\frac{9\times10^9\times2\times5\times10^{-18}}{0.02^2}=2.25\times10^{-4}N"

"F_3=\frac{kq_3q_2}{r^2}"

"F_3=-\frac{9\times10^9\times3\times5\times10^{-18}}{0.04^2}=-8.4375\times10^{-5}N"

"F_{net}=F_1+F_3"


"F_{net}=2.25\times10^{-4}-8.4375\times10^{-5}=1.406\times10^{-4}N"